Marky:

1. The length of the rectangular table is 4 feet longer than its width. Its area is 12 square feet. What is its length?

Marky:

Wdym? Choice?

Marky:

No it didn't have any choices that's why i need your help

supie:

@mxddi3 can you help?

Marky:

And deadline is today

Marky:

Help

jimthompson5910:

x = width x+4 = length, since the length is 4 ft longer than the width The length and width multiply to the area of 12 sq ft (length)*(width) = area of rectangle (x+4)*x = 12 x^2+4x = 12 x^2+4x-12 = 0 What's the next step from here?